# Super approximation for $\text{SL}_{2}\times\text{SL}_{2}$ and $\text{ASL}_{2}$

### v3

@@ 4. Preliminaries on Random Walks
=\ = $\chi_{S}^{(l)}(\mathcal{G})$
[…]
<\ < $e^{-c_{0}l}.$
where $c_{0}$ is the constant given in Lemma 4.7.
Now let us deal with the case $n\neq 0$. Let $C_{3}=3^{70}\mathfrak{X}(8)$, where $\mathfrak{X}(8)$ is the constant given by Theorem C, and take $c_{S}=\frac{1}{C_{3}}$. Write $Q_{1}=(Q,n)$.
Suppose $Q_{1}<Q^{\frac{1}{C_{3}}}$. First we can find $n^{\prime}=Q_{1}m\equiv 0(\operatorname{mod}Q)$ with $(m,Q)=1$. The equation $L(\gamma)=n(\operatorname{mod}Q)$ is the same as $L(\gamma)=Q_{1}m(\operatorname{mod}Q)$. Since $m$ is invertible mod $Q$, we can define
$f_{\gamma}(X,Y,Z,W,A,B,C,D)=xX+yY+zZ+wW+aA+bB+cD+dD-Q_{1}$
for $\gamma=\left(\left(\begin{matrix}x&y\\ z&w\end{matrix}\right),\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)$, and proceed as the $n=0$ case. The reason in the $n=0$ case we do not define $f_{\gamma}$ in this way is that we need to ensure a solution to $f_{\gamma}=0$ is nonzero. But if $n\neq 0$, $f_{\gamma}=0$ never admits a zero solution.
If $Q_{1}\geq Q^{\frac{1}{C_{3}}}$, then we can reduce this case to the $n=0$ case by considering the equation $L(g)\equiv 0(\operatorname{mod}Q_{1})$ instead of $L(g)\equiv n(\operatorname{mod}Q)$, and letting $c_{S}=\frac{1}{8C_{2}C_{3}}$. ∎
5. Bounded generation for simple factors
Let $\mathbb{P}_{0}$ be the projection from $\text{SL}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}$ to $\text{SL}_{2}(\mathbb{Z})$, and let $\mathbb{P}_{1}$ ($\mathbb{P}_{2}$, resp.) be the projection map from $\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ to its first factor (second factor, resp.). The goal of this section is to prove, the following bounded generation result for simple factors of $\Lambda(\operatorname{mod}q)$, where $q$ is a product of large powers of primes:
Proposition 5.1.
Let $S,\epsilon$ be as given in Proposition 2.2. There are constants $c>0$ depending only on the generating set $S$ and $\epsilon$, and a constant $\rho=\rho(\delta)>0$, $C=C(\delta)\in\mathbb{Z}_{+}$, with $\rho=O_{S}(\delta)$, such that for any $q=\prod_{i\in I}p_{i}^{n_{i}},n_{i}>\frac{12}{\delta}$, if the conclusion of Proposition 2.2 fails, i.e.,
$|\pi_{q}(A\cdot A\cdot A)|\leq|\pi_{q}(A)|^{1+\delta},$ (5.2)
then we have
$\Gamma(q_{*}^{\{\rho\}})/\Gamma(q_{*})\subset\begin{cases}&(\mathbb{P}_{i}A)^{C}(\operatorname{mod}q_{*}),i=1,2,\hskip 8.53581pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\times{\text{SL}}_{2}(\mathbb{Z}),\\ &\mathbb{(}P_{0}A)^{C}(\operatorname{mod}q_{*}),\hskip 45.5244pt\text{if }\hskip 5.69054pt\Lambda={\text{SL}}_{2}(\mathbb{Z})\ltimes\mathbb{Z}^{2}.\end{cases}$ (5.3)
where $q_{*}||q$ and $q_{*}\geq q^{c}$.
We focus on the case $\Lambda=\text{SL}_{2}(\mathbb{Z})\times\text{SL}_{2}(\mathbb{Z})$ and $i=1$. The proofs for the other two cases are identical. Our proof mostly follows the steps of [BG09].
5.1. A large set of commuting elements
We use Helfgott’s argument [Hel08] to create a large set of commuting elements under an appropriate modulus.
Let $c_{1},c_{2}$ be the implied constants for $c$ from Proposition 4.1 and Proposition 4.2 respectively. Let $A_{0}=A\cdot A\cap\Lambda(q_{0})$, where $q_{0}=\prod_{p|q_{l}}p$. Then
$\chi_{S}^{2l}(A_{0})>q^{-2\delta}q_{0}^{-12}>q^{-3\delta},$ (5.4)
where we have used the exponent for each prime divisor of $q$ exceeds $\frac{12}{\delta}$.
Let $\nu$ be the push forward of $\chi_{S}$ under the projection map $\mathbb{P}_{1}$ and let
$A_{1}=\mathbb{P}_{1}(A_{0}).$
For each $Q|q$, let
$\mathcal{E}_{1}(Q)=\{\gamma\in\text{SL}_{2}(\mathbb{Z}):\text{tr}(\gamma)^{2}-4\equiv 0(\operatorname{mod}Q))\},$
Then Proposition 4.1 gives
$\nu^{(2l)}(\mathcal{E}_{1}(Q)),\nu^{(2l)}(\mathcal{E}_{2}(Q))<Q^{-c_{1}}.$
Take $\rho_{1}>0$, and let
$A_{2}=A_{1}-\cup_{\begin{subarray}{c}Q|q\\ Q\geq q^{\{\rho_{1}\}}\end{subarray}}(\mathcal{E}_{1}(Q)\cup\mathcal{E}_{2}(Q)),$
We should think of $\rho_{1}$ as a relative level depending only on $S$, so as $\rho_{2}$ in the following; the exact value of $\rho_{1}$ is determined at (5.58). Since the number of divisors of $q$ is bounded by $q^{0+}$, it follows that
$\nu^{(2l)}(A_{2})>q^{-3\delta-},$
if we let
$\delta<\frac{\rho_{1}c_{1}}{4}.$ (5.5)
A simple argument by contradiction shows that for each $\gamma\in A_{2}$, there exists $q_{\gamma}||q$, $q_{\gamma}>q^{\frac{1}{2}}$, such that for any $p^{n}||q_{\gamma}$,
$\text{tr}(\gamma)^{2}-4\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}),$ (5.6)
$\gamma_{21}\neq 0(\operatorname{mod}p^{[4\rho_{1}n]}).$ (5.7)
Since again the number of divisors of $q$ is bounded by $q^{0+}$, there exists $q_{1}||q,q_{1}>q^{\frac{1}{2}}$ and a set $A_{3}\subset A_{2}$, with
$\nu^{(2l)}(A_{3})>q^{-3\delta-},$ (5.8)
[…]
@@ 5.1. A large set of commuting elements
$Q_{1}=q_{1}^{\{\rho_{2}\}},$
where we will take $\rho_{2}\gg\rho_{1}$.

### v4

@@ 5.2. A large set of commutative matrices
which contradicts (5.22) for $l=4$, if we take $L_{2}$ sufficiently large.
For each $\alpha\in A_{1}$, let $C_{\alpha}=\{\gamma\alpha\gamma^{-1}:\gamma\in\mathbb{P}_{1}(A)\}$. Since distinct elements in $A_{1}$ have distinct eigenvalues (mod $Q_{0}$), the sets $\{\pi_{Q_{0}}(C_{\alpha})\}_{\alpha\in A_{1}}$ are mutually disjoint. Therefore,
$\sum_{\alpha\in A_{1}}|\pi_{Q_{0}}(C_{\alpha})|=|\cup_{\alpha\in A_{1}}\pi_{Q_{0}}(C_{\alpha})|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A)^{4})|<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{18}}.$
Thus, recalling (5.8), there exists some $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)\in A_{1}$,
$|\pi_{Q_{0}}(C_{\alpha_{0}})|\leq|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|^{1+\frac{c_{1}}{18}}Q_{0}^{-c_{1}/2}<|\pi_{Q_{0}}(\mathbb{P}_{1}(A))|Q_{0}^{-\frac{c_{1}}{3}}.$ (5.26)
It follows from (5.26) that there exists $x_{0}\in\mathbb{P}_{1}(A)$ and
$A_{2}^{\prime}:=\{x\in\mathbb{P}_{1}(A):x\alpha_{0}x^{-1}=x_{0}\alpha_{0}x_{0}^{-1}(\operatorname{mod}Q_{0})\},$ (5.27)
such that
$|\pi_{Q_{0}}(A_{2}^{\prime})|>Q_{0}^{\frac{c_{1}}{3}}.$ (5.28)
Therefore, $A_{2}^{\prime\prime}=A_{2}^{\prime}\cdot x_{0}^{-1}$ commutes with $\alpha_{0}$ mod $Q_{0}$ and
$\displaystyle|\pi_{Q_{0}}(A_{2}^{\prime\prime})|>Q_{0}^{\frac{c_{1}}{3}},$ (5.29)
Choose $A_{2}\subset A_{2}^{\prime\prime}$ to be a set of representatives for $\pi_{Q_{0}}(A_{2}^{\prime\prime})$, which gives Lemma 5.18. ∎
5.3. Applying sum-product
To proceed, we first diagonalize $\alpha_{0}$ in an extended number field. Let
be the two eigenvalues of $\alpha_{0}$. Let $K=\mathbb{Q}[\lambda_{1}]$, and $\mathcal{O}=\mathcal{O}_{K}$ be the ring of integers of $K$. Let $M=\left(\begin{matrix}\lambda_{1}-\alpha_{22}&\lambda_{2}-\alpha_{22}\\ \alpha_{21}&\alpha_{21}\end{matrix}\right)$. Then
$M^{-1}\alpha_{0}M=\left(\begin{matrix}\lambda_{1}&0\\ 0&\lambda_{2}\end{matrix}\right).$ (5.30)
Since $\alpha_{0}$ satisfies the condition (5.10) for $\gamma$,
$\lambda_{1}\not\equiv\lambda_{2}(\operatorname{mod}p^{[4\rho_{1}n]}),\hskip 28.45274pt\forall p^{n}\|q_{1}.$ (5.31)
Let $C_{1},C_{2},C_{3},\varepsilon=\varepsilon_{0}$ be the implied constants of Theorem B for $\alpha=\frac{c_{1}}{6}$.
The first goal of this section is to prove
Lemma 5.32.
There are scales $0<w_{1},w_{2}<1$ with $w_{1},w_{2}=\Theta(\rho_{2})$, $w_{2}\leq w_{1}$, and a set $H\subset M^{-1}\cdot A_{2}^{2}\cdot M$, and an ideal $\mathcal{H}\supset\left(q_{1}^{\{\frac{w_{2}}{C_{1}}\}}\right)$ in $\mathcal{O}$, such that
  1. (i)
    $H\equiv 1(\operatorname{mod}q_{1}^{\{w_{1}\}})$,
  2. (ii)
    $H$ is diagonal mod $q_{1}^{\{w_{1}+w_{2}\}}$,
  3. (iii)
    (lower bound for dimension) Elements in $H$ are distinct mod $q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}$ and
    $\left|\pi_{q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}}(H)\right|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{5}-}.$
  4. (iv)
    (common divisor) For each $h\in H$, writing $h\equiv\left(\begin{matrix}a_{h}&0\\ 0&a_{h}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}})$ for some $a_{h}\in\mathcal{O}$, then $\langle{a_{h}-1},q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}\rangle=q_{1}^{\{w_{1}\}}\mathcal{H}$.
Remark 5.33.
The reason that we choose the scales $w_{1},w_{2}$ in the above way is that subsequently we will construct a set $\mathscr{H}\subset\mathcal{O}$ from $H$ and apply Theorem B to $A=\mathscr{H}$, $\mathfrak{a}=(q_{1}^{\{\frac{w_{2}}{C_{1}}\}})$ and density $\alpha=\frac{c_{1}}{6}$. The conclusion of Theorem B then ensures that the obtained ideal $\mathfrak{a}^{\prime}\supset(q_{1}^{\{w_{2}\}})$, so that we can use Property (ii) from Lemma 5.32 in the computation of some matrix products for the proof of Lemma 5.53.
Proof of Lemma 5.32.
The denominator of $M^{-1}$ is controlled $\text{Det}(M)=\alpha_{21}\sqrt{\text{tr}(\alpha_{0})^{2}-4}$, from which together with (5.10), we have
$p^{[8\rho_{1}n]}\nmid\text{Det}(M),\hskip 28.45274pt\forall p^{n}\|q_{1}.$ (5.34)
Therefore, from Lemma 5.18, since $A_{2}$ commutes with $\alpha_{0}$ mod $q_{1}^{\{\rho_{2}\}}$, we have
$H_{0}:=M^{-1}A_{2}M$
commutes with $M^{-1}\alpha_{0}M$ mod $q_{1}^{\{\rho_{2}-8\rho_{1}\}}$. Moreover, because of the condition (5.31), we have, for any $\gamma\in A_{2}$, we can find $a_{\gamma}\in\mathcal{O}$ such that
$M^{-1}\gamma M\equiv\left(\begin{matrix}a_{\gamma}&0\\ 0&a_{\gamma}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}}).$ (5.35)
We first give an estimate for $|H_{0}|$. The density condition (5.19) implies that
$|A_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}-20\rho_{1}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{4}},$ (5.36)
by taking $L_{2}$ sufficiently large.
If $\gamma_{1}\not\equiv\gamma_{2}(\operatorname{mod}q_{1}^{\{\rho_{2}-20\rho_{1}\}})$, then $M^{-1}\gamma_{1}M\not\equiv M^{-1}\gamma_{2}M(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})$. (5.36) thus implies:
$|H_{0}(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})|>q_{1}^{\frac{\rho_{2}c_{1}}{4}}.$ (5.37)
We should think of the $\rho_{1}$ terms in (5.35), (5.36), (5.37) negligible compared to the $\rho_{2}$ terms, and this applies to any similar expressions.
The density condition (5.37) implies that at east one of the following two events occurs: Event 1: There is a scale $w_{1}\in[\frac{c_{1}}{80}\rho_{2},\frac{1}{2}(\rho_{2}-12\rho_{1})]$, and $h_{0}\in H_{0}$ such that the set
$H_{1}:\{h\in H_{0}:h\equiv h_{0}(\operatorname{mod}q_{1}^{\{w_{1}\}})\}$ (5.38)
satisfies
$|H_{1}(\operatorname{mod}q_{1}^{\{w_{1}(1+C_{1}^{-1})\}})|>q_{1}^{\frac{w_{1}}{C_{1}}\cdot\frac{c_{1}}{5}},$ (5.39)
In this case, we set $w_{2}=w_{1}$.
Event 2: There is a scale $w_{1}\in[\frac{1}{2}(\rho_{2}-12\rho_{1}),(1-\frac{c_{1}}{80})\rho_{2}]$ and $h_{0}\in H_{0}$ such that the set
$H_{1}:\{h\in H_{0}:h\equiv h_{0}(\operatorname{mod}q_{1}^{\{w_{1}\}})\}$ (5.40)
satisfies
$|H_{1}(\operatorname{mod}q_{1}^{\{w_{1}+\frac{\rho_{2}-12\rho_{1}-w_{1}}{C_{1}}\}})|>q_{1}^{\frac{\rho_{2}-12\rho_{1}-w_{1}}{C_{1}}\cdot\frac{c_{1}}{5}},$ (5.41)
In this case, we set $w_{2}=\rho_{2}-12\rho_{1}-w_{1}$.
Indeed, suppose no such $w_{1}$ and $h_{0}$ exist. Set $t_{0}=\frac{c_{1}}{80}\rho_{2},$ and $t_{s}=t_{0}(1+C_{1}^{-1})^{s}$ for $1\leq s\leq s_{1}$, where $s_{1}$ is the largest integer such that $t_{s_{1}}\leq\frac{1}{2}(\rho_{2}-12\rho_{1})$. We continue to iteratively define $t_{s_{1}+1}=t_{s_{1}}+\frac{\rho_{2}-12\rho_{1}-t_{s_{1}}}{C_{1}}$, $t_{s_{1}+2}=t_{s_{1}+1}+\frac{\rho_{2}-12\rho_{1}-t_{s_{1}+1}}{C_{1}}\cdots$, until we reach a first integer $s_{2}$ such that $t_{s_{2}}>(1-\frac{c_{1}}{80})\rho_{2}$. Then
$|H_{0}(\operatorname{mod}q_{1}^{\{\rho_{2}-12\rho_{1}\}})|$ $\leq|\pi_{q_{1}^{\{t_{s_{2}}\}}}(H_{0})|\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
$\leq|\pi_{q_{1}^{\{t_{s_{2}-1}\}}}(H_{0})|\cdot q_{1}^{({t_{s_{2}}-t_{s_{2}-1}})\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
$\leq|\pi_{q_{1}^{\{t_{s_{2}-2}\}}}(H_{0})|\cdot q_{1}^{({t_{s_{2}-1}-t_{s_{2}-2}})\cdot\frac{c_{1}}{5}}\cdot q_{1}^{({t_{s_{2}}-t_{s_{2}-1})}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
$\cdots$
$\leq|\pi_{q_{1}^{\{t_{0}\}}}(H_{0})|\cdot q_{1}^{{t_{s_{2}}}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}$
$\leq q_{1}^{\frac{\rho_{2}c_{1}}{40}}\cdot q_{1}^{{t_{s_{2}}}\cdot\frac{c_{1}}{5}}\cdot q_{1}^{\frac{\rho_{2}c_{1}}{40}}\leq q_{1}^{\frac{\rho_{2}c_{1}}{4}},$
contradicting (5.37). So such $h_{0}$ and $w_{1}$ have to exist.
Let $H_{2}\subset h_{0}^{-1}H_{1}$ be a set of representatives of $h_{0}^{-1}H_{1}(\operatorname{mod}q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}})$. Then $H_{2}$ satisfies the conditions (i), (ii), (iii) for Lemma 5.32.
For each $h\in h_{0}^{-1}H_{2}$, take some $a_{h}\in\mathcal{O}$ so that
$h\equiv\left(\begin{matrix}a_{h}&0\\ 0&a_{h}^{-1}\end{matrix}\right)(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}}).$
From (5.40), for each $h\in H_{2}$, we have $a_{h}\equiv 1(\operatorname{mod}q_{1}^{\{w_{1}\}})$. Let $\mathcal{H}_{h}$ be an ideal of $\mathcal{O}$ such that $q_{1}^{\{w_{1}\}}\mathcal{H}_{h}=\langle a_{h}-1,q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}}\rangle$. Since the number of divisors of $\left(q_{1}^{\left\{\frac{w_{2}}{C_{1}}\right\}}\right)$ is $O(q_{1}^{0+})$, there is $\mathcal{H}$ such that
$H:=\{h\in H_{2}:\mathcal{H}_{h}=\mathcal{H}\}$
satisfies
$|H(\operatorname{mod}q_{1}^{\{w_{1}+\frac{w_{2}}{C_{1}}\}})|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{5}-}.$ (5.42)
Then $H$ satisfies all requirements in Lemma 5.32.
For each $h\in H$, there is $\lambda_{h}\in\mathcal{O}$ such that
$a_{h}-a_{h}^{-1}\equiv q_{1}^{\{w_{1}\}}\lambda_{h}(\operatorname{mod}q_{1}^{\{w_{1}+w_{2}\}}).$ (5.43)
@@ 5.3. Applying sum-product
$\mathscr{H}:=\{\lambda_{h}:h\in H_{2}\}.$
From (5.42) and the fact that for an arbitrary ideal $\mathfrak{a}\subset\mathcal{O}$, the map
$(\mathcal{O}/\mathfrak{a})^{*}\rightarrow\mathcal{O}/\mathfrak{a}:x\rightarrow x-x^{-1}$
is $O(|\mathcal{O}/\mathfrak{a}|^{0+})$ to one, we have
$|\mathscr{H}(\operatorname{mod}q_{1}^{\{\frac{w_{2}}{C_{1}}\}})|>q_{1}^{\frac{w_{2}}{C_{1}}\cdot\frac{c_{1}}{6}}.$ (5.44)
For a construction in the next section, we apply Theorem B to the set $\mathscr{H}$, the density $\alpha=\frac{c_{1}}{6}$ and the ideal $\mathfrak{a}=(q_{1}^{\{\frac{w_{2}}{C_{1}}\}})$, which gives us an ideal $\mathcal{Q}\subset\mathfrak{a}^{C_{1}}=(q_{1}^{\{w_{2}\}})$, $\xi\in\mathcal{O}$, and some $\varepsilon_{0}>0$ depending only on $c_{1}$, such that
$\pi_{\mathcal{Q}}(\mathbb{Z}\xi)\subset\pi_{\mathcal{Q}}\left(\sum_{C_{3}}\mathscr{H}^{C_{2}}-\sum_{C_{3}}\mathscr{H}^{C_{2}}\right),$ (5.45)
$|\pi_{\mathcal{Q}}(\mathbb{Z}\xi)|>q_{1}^{{w_{2}\varepsilon_{0}}}.$ (5.46)
Remark 5.47.
In fact, following the proof of Theorem B, the ideal $\mathcal{Q}$ satisfies the property that different prime ideals $\mathcal{P}|\mathcal{Q}$ live above different natural primes $p$. We will assume this property of $\mathcal{Q}$ in the following discussion.
5.4. Constructing a line
Let $q_{2}$ be the product of all $p^{n}\|q_{1}$ where the projection of the arithmetic progression (5.46) localized at $p$ is large, i.e. write $\mathcal{Q}_{p}=\langle\mathcal{Q},p^{n}\rangle$, and let
$q_{2}:=\prod_{\begin{subarray}{c}p^{n}\|q_{1}\\ |\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|>p^{nw_{2}\cdot\frac{\varepsilon_{0}}{2}}\end{subarray}}p^{n}.$ (5.48)
The density condition (5.46) implies
$q_{2}>q_{1}^{{\varepsilon_{0}}/{4}}.$ (5.49)
Indeed, suppose not, then
$|\pi_{\mathcal{Q}}(\mathbb{Z}\xi)|=\prod_{p^{n}\|q_{2}}|\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|\cdot\prod_{p^{n}\|\frac{q_{1}}{q_{2}}}|\pi_{\mathcal{Q}_{p}}(\mathbb{Z}\xi)|\leq q_{2}^{2}(q_{1}/q_{2})^{\frac{w_{2}\varepsilon_{0}}{2}}\leq q_{1}^{w_{2}\varepsilon_{0}},$
contradicting (5.46).
For each $p|q_{2}$, let $\mathcal{P}_{p}|\mathcal{Q}$ be the unique prime ideal that lives above $p$ (see Remark 5.47), and $\mathcal{Q}_{p}$ be the $\mathcal{P}_{p}$ power that exactly divides $\mathcal{Q}$. In our case, the extension degree is at most 2, so $v_{\mathcal{P}_{p}}(p)=1$ or 2.
The following is our goal in this section:
Lemma 5.50.
There is a positive integer $K_{1}$ depending only on $S$, an exact divisor $q_{4}\|q_{2}$, $q_{4}\geq q_{2}^{1/4}$, rational integers $Q_{1}|Q_{2}|q_{4}$, a traceless matrix $X=\left(\begin{matrix}a&b\\ c&-a\end{matrix}\right)\in{\text{Mat}}_{2}(\mathbb{Z})$ coprime to $q_{4}$, such that
$1+\mathbb{Z}Q_{1}X\subset A_{0}^{K_{1}}(\operatorname{mod}Q_{2}),$
where for each $p^{n}\|q_{4}$, we have
$v_{p}(Q_{2})=\Theta(\rho_{2}n)$ (5.51)
$v_{p}(Q_{2})-v_{p}(Q_{1})=\Theta(\rho_{2}n).$ (5.52)
We first use random walk to find an element from a product of $A_{0}$ which is very non-commutative with $\alpha_{0}=\left(\begin{matrix}\alpha_{11}&\alpha_{12}\\ \alpha_{21}&\alpha_{22}\end{matrix}\right)$ mod $q_{2}$, where $\alpha_{0}$ is given as in Lemma 5.18, and then conjugate it by $M$ so as to obtain a non-diagonal matrix. The following is our goal:
Lemma 5.53.
There is an exact divisor $q_{4}\|q_{2}$ with $q_{4}\geq q_{2}^{\frac{1}{4}}$, an ideal $\mathscr{Q}_{1}\supset(q_{4})$, an element $\zeta\in M^{-1}A_{0}HA_{0}^{-1}H^{-1}M$, such that
$\zeta\equiv 1+\mathfrak{q}_{1}\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)(\operatorname{mod}\mathscr{Q}_{1}^{2})$
where $\left(\begin{matrix}X&Y\\ Z&-X\end{matrix}\right)$ is a primitive matrix in $\text{Mat}_{2}(\mathcal{O})$, and $\mathfrak{q}_{1}$ is an element in $\mathcal{O}$ that uniformizes $\mathscr{Q}_{1}$. For each $p^{n}\|q_{4}$, we have the following estimates for valuations:
$0.99v_{\mathcal{P}_{p}}(p^{[nw_{1}]})\leq v_{\mathcal{P}_{p}}(\mathscr{Q}_{1})\leq 1.01v_{\mathcal{P}_{p}}(p^{[nw_{1}]})$ (5.54)
where $w_{1}$ is given as in Lemma 5.32 and
$v_{\mathcal{P}_{p}}(Z)=O(n\rho_{1}).$ (5.55)
Proof.
We work with the linear form
$\mathcal{L}\left(\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\right)=\alpha_{21}a+(\alpha_{22}-\alpha_{11})c-\alpha_{21}d,$ (5.56)
which gives the 2-1 entry of $\alpha_{0}\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)-\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)\alpha_{0}$. If (5.56) is non-zero, this certainly implies $\zeta$ and $\left(\begin{matrix}a&b\\ c&d\end{matrix}\right)$ are not commutative.